# ResearchBot: Bridging the Gap between Academic Research and Practical Programming Communities

### v1

@@ Abstract.
Software developers commonly rely on platforms like Stack Overflow for problem-solving and learning. However, academic research is an untapped resource that could greatly benefit industry practitioners. The challenge lies in connecting the innovative insights from academia to real-world problems faced by developers. This project introduces ResearchBot, a tool designed to bridge this academia-industry gap. ResearchBot employs a modular approach, encompassing understanding questions, curating queries to obtain relevant papers in the CrossRef repository, summarizing paper content and finally answering user questions based on paper summaries. The core objective of ResearchBot is to democratize access to academic knowledge for industry professionals. By providing concise summaries of cutting-edge research directly in response to SE-related questions, ResearchBot facilitates the application of academic insights to practical contexts. Ultimately, it aims to bridge the gap between academia and industry, using research evidence to support learning and decision-making in software development.
Demo URL: https://youtu.be/ny6TIoNAor8
GitHub: https://github.com/Sfrznp/ResearchBot
Data Set: https://tinyurl.com/y3cu9mvx
Software Engineering, Software Tools, Natural Language Processing (NLP)
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 17–01, 2024; Sacramento, CAisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software notations and tools
1. Introduction
In software engineering (SE), there is a significant gap between academic research and practical programming communities. This gap is caused by challenges such as the accessibility, relevance, and credibility of research when applied to the fast-paced field of software development (Winters, 2024). Despite the rigorous methods and extensive analyses used in SE research, evidence from scientific insights often fail to reach practitioners seeking solutions. Community-driven platforms (e.g., Stack Overflow [^1] https://stackoverflow.com/) are frequently used to support learning and problem-solving in SE (Barua et al., 2014). More recently, large language models (LLMs, e.g., ChatGPT [^2] https://chat.openai.com/) are capable of providing solutions to SE-related tasks (Leung and Murphy, 2023). However, prior work highlights issues regarding academic findings in Stack Overflow (Huang and Chattopadhyay, 2024) and ChatGPT (Auer et al., 2023), emphasizing the need for novel approaches to align research evidence with practical community needs.
@@ 1. Introduction
To address these challenges, we introduce ResearchBot, an early prototype of a chatbot designed to make academic research more accessible and relevant to programming practitioners. ResearchBot leverages advanced natural language processing (NLP) techniques to automatically understand user questions, retrieve relevant papers from the CrossRef repository [^3] https://www.crossref.org/—a comprehensive database containing metadata about scholarly articles, conference proceedings, books, and datasets, and provide concise summaries. Our preliminary evaluation explores the capabilities of ResearchBot to provide evidence-based and human-like responses to practical SE-related queries. By transforming theoretical knowledge into practical contexts, ResearchBot aims to provide a first step in bridging the gap between academia and industry—facilitating the application of research evidence to real-world SE problems (Kitchenham et al., 2004).
2. Implementation
Figure 1. Pipeline for the implementation of ResearchBot
@@ 2. Implementation
The implementation of ResearchBot integrates two main modules: the Retrieval Module and the Question-Answer (QA) Module, leveraging GPT-4’s superior capabilities of natural language understanding. This system utilizes specific prompts to guide the model in querying the Crossref repository and generating informative responses based on retrieved research papers. Figure 1 illustrates the pipeline of ResearchBot, showcasing the flow from query to response generation.
2.1. Retrieval Module
@@
The Retrieval Module of ResearchBot is responsible for the process of retrieving relevant academic papers based on user queries.
2.1.1. Prompt used to instruct GPT-4 to curate a query for the Crossref API:
From now on I will pass in a prompt like "How do I
perform document similarity using NLP" and I would
like you to  curate an output that looks similar to
this "nlp+natural+language+processing+document+
similarity". This is what a search query would look
like for the given prompt. It uses ’+’ which
separates all the key words.
2.1.2. Significance of Prompt:
The prompt used in this module is beneficial because it guides GPT-4 on how to transform a natural language question into a structured query format. This ensures that the search query generated by the model is specific and optimized for retrieving relevant academic papers. Without a clear prompt, the model might struggle to understand the intended format of the query and could produce less accurate results.
2.1.3. Workflow:
  1. (1)
@@ 2.1.3. Workflow:
    (e.g., ”nlp+natural+language+processing+document+
@@
2.2. QA Module
@@
2.2.1. Prompt used to instruct GPT-4 to answer the user’s question:
From now on I will pass in a Software Engineering
question and some relevant research papers as
context for the answer. I would like you to answer
the question with references made to the contexts
(make sure that the focus of the answer is to
answer the question, not the context itself).
Be sure to also include the URL after the reference
to the paper in the text. Use this format to cite
in-text (<a href="url" target="_blank">title</a>).
2.2.2. Significance of Prompt:
The prompt used in the QA Module is crucial as it provides GPT-4 with clear instructions on how to structure its response. By specifying the focus on answering the user question with reference to the provided context (research paper abstracts), the prompt ensures that the generated response is relevant and informative. Additionally, the prompt dictates the citation format, enabling proper acknowledgment of the cited sources within the response.
2.2.3. Workflow:
  1. (1)
@@ 2.2.3. Workflow:
2.3. Reasoning behind the use of LLMs
LLMs powered by generative artificial intelligence (AI), such as GPT-4, offer several advantages for ResearchBot:
  • Natural Language Understanding: LLMs excel in understanding and generating human-like text, enabling them to process complex queries and produce coherent responses.
  • Contextual Awareness: GPT-4 can leverage context from multiple sources (e.g., abstracts of research papers) to provide informed answers tailored to specific questions.
  • Flexibility and Adaptability: The model’s flexibility allows it to adapt to various domains and tasks, making it suitable for bridging academic literature with practical programming challenges.
3. Methodology
3.1. Data Collection
@@
For the purpose of this study, data was systematically collected from the Software Engineering Stack Exchange community. [^4] https://softwareengineering.stackexchange.com/ This platform was selected due to its extensive use by programming professionals, who regularly engage in discussions concerning practical challenges and their solutions. Our selection criteria focused primarily on technical inquiries, particularly those pertaining to subjects such as programming languages and software development. Recognizing the potential limitation in our bot’s ability to accurately respond to coding-related questions, we avoided using inquiries from Stack Overflow. However, we still included some coding questions from SE Stack Exchange in our evaluation. Using these criteria, we collected 100 questions to analyse the functionality of ResearchBot.
3.2. Data Analysis
3.2.1. Semantic Textual Similarity
@@
We utilized semantic textual similarity (STS), an NLP technique for determining how similar two pieces of text are in terms of meaning, to test and analyze our ResearchBot’s performance and accuracy. The primary goal of this analysis is to compare the likeness in meaning between the user’s answer to a question on SE Stack Exchange and ResearchBot’s generated response. The higher their similarity, the better our tool answers the user’s question. To achieve our goal, we used the Sentence-BERT  [^5] https://sbert.net/ (SBERT) model and the similarity method from spaCy  [^6] https://spacy.io/usage/linguistic-features#vectors-similarity, an open-source NLP library in Python.
3.2.2. LLM Comparison
To further assess ResearchBot’s performance, we compared its responses with those generated by other LLM tools commonly used in SE: Google Gemini, GitHub Copilot, and ChatGPT. This evaluation involved 50 top questions from Stack Exchange. Upon manual review of the answers, we found responses from Gemini and Copilot were unreliable and lacked quality references. ChatGPT offered better resources; thus, we use ChatGPT as the state-of-the art platform to compare against ResearchBot.
3.2.3. Analysis Process
@@
Firstly, we selected 100 question-and-answer sets for testing and analysis. We then wrote a script to generate corresponding answer outputs for the 100 questions by ResearchBot and ChatGPT. Next, we used the STS tools, SBERT and spaCy, to compare the answers most upvoted from users (top answer) and answers produced by ResearchBot (generated answer). We use the top answer to represent the correct answer, most accepted by humans with similar inquiries. Finally, we calculated each comparison set’s semantic similarity score and each STS tool’s overall statistic measurement characteristics of data. We also employed the STS tools to compare generated answers with those produced by ChatGPT. This analysis may introduce biases, as it is based on the assumption that the top answer is both correct and precise. This could lead to overestimating the bot’s performance, as similarity with the top answer might not represent the most accurate or comprehensive response. To enhance the validity of our results, the incorporation of a broader range of benchmarks and metrics are recommended.
4. Results
@@
This section will present the results of our data collection and analysis approach. Table 1 and Table 2 present the statistical results for semantic similarity comparing the performance of ResearchBot with Stack Exchange and ChatGPT, respectively.
Table 1. Stack Exchange vs. ResearchBot
@@ 4.1. STS Results
The results presented in Table 1 demonstrate that ResearchBot performs effectively in answering users’ questions as analyzed by spaCy, with an average similarity score of 0.92. However, the performance analyzed by SBERT shows a lower average of 0.66. This discrepancy is largely due to a notably low minimum similarity score of 0.0341 recorded by SBERT. This result highlights SBERT’s higher scrutiny evaluation criteria, contributing to higher variability in scores, as indicated by a standard deviation of 0.16, compared to spaCy. Such variability suggests a re-evaluation of spaCy’s data handling, as it indicates methodological differences, suggesting a need for a deeper comparative analysis of how each tool handles and interprets textual variations. Further analysis could provide insights into the different sensitivities of the tools to specific text features or contexts, thus explaining the variations in their performance metrics. Moreover, similar to recent efforts in sentiment analysis, future work can explore SE-specific NLP tools for STS.
@@
Overall, the answer outputs provide further insights, demonstrating ResearchBot not only matches the top answers but also provides more comprehensive and detailed responses with research-based evidence. This ability to elaborate and expand on critical points makes the tool particularly valuable. Consequently, while ResearchBot shows promise, there is room for improvements of our tool and evaluation in answering SE-related questions.
4.2. Comparison Results
We compared ResearchBot with state-of-the-art ChatGPT in responses to SE-related questions. We employed the STS tools to compare the answers generated by ChatGPT and ResearchBot. The scores are detailed in Table 2. ResearchBot achieved an average score of 0.9525 with spaCy and 0.8175 with SBERT, indicating that its answers are quite similar to those from ChatGPT. However, a closer examination revealed issues with the evidence provided in ChatGPT responses. It is worth noting that, though some answers did reference academic papers, we observed several instances of broken links and unreliable websites. ResearchBot, in contrast, demonstrated superior reliability and resource quality in its answers. To further assess the responses, we are conducting a user study to determine user preferences between ChatGPT and ResearchBot.
Table 2. ChatGPT vs. ResearchBot
Sentence-BERT model 0.9323 0.7137 0.8175 0.8195 0.0426
spaCy library 0.9874 0.8432 0.9525 0.9666 0.0506
5. Discussion
5.1. Implication and Impact of the Results
@@
By utilizing STS tools and comparing against mainstream LLMs, we gained valuable insights into how ResearchBot performs in generating accurate and relevant responses to user inquiries. The results show a mean STS scores of 0.66 (SBERT) and 0.92 (spaCy), indicating accurate human-like responses to practical SE-related questions. Further manual analysis shows that, despite lower scores for SBERT, ResearchBot’s answers are not only correct but provide further background to answer the user’s question. We also show our system responses are comparable to ChatGPT while providing more reliable responses based on research evidence. The implications of our study extend to practical contexts, motivating future tools to facilitate theoretical knowledge (as in non-coding questions—i.e., conceptual questions) transfer between academia and industry. By enhancing ResearchBot’s ability to provide accurate and informative responses based on academic knowledge, we contribute to bridging the academic-industrial gap in SE.
5.2. Future Work
@@
User Study to Further Evaluate the Responses.
A limitation of our preliminary study is we rely on STS comparisons with top answers from SE Stack Exchange and ChatGPT responses to evaluate the capabilities of ResearchBot. As previously mentioned, we are currently conducting a user study where participants are asked to choose between answers to practical SE-related questions generated by ResearchBot and ChatGPT, providing a rationale for their preference. This study aims to provide a better understanding of the types of answers users prefer, motivate enhancements to ResearchBot, and further promote evidence-based practices in SE contexts.
Expansion of Academic Repositories.
@@
User Interaction and Feedback Mechanisms.
@@
Code Understanding, Creation, and Correction.
@@
Integration with Programming Platforms.
We also envision exploring integration opportunities for our system in popular programming platforms beyond SE Stack Exchange, such as GitHub or developer forums. This can expand ResearchBot’s reach and utility within the broader SE community. Further, integration into integrated development environments (IDEs) can also facilitate insights from ResearchBot during software development to answer questions and inform decision making.
6. Conclusion

### v2

@@ Abstract
Software developers commonly rely on platforms like Stack Overflow for problem-solving and learning. However, academic research is an untapped resource that could greatly benefit industry practitioners. The challenge lies in connecting the innovative insights from academia to real-world problems faced by developers. This project introduces ResearchBot, a tool designed to bridge this academia-industry gap. ResearchBot employs a modular approach, encompassing understanding questions, curating queries to obtain relevant papers in the CrossRef repository, summarizing paper content, and finally answering user questions based on paper summaries. The core objective of ResearchBot is to democratize access to academic knowledge for industry professionals. By providing concise summaries of cutting-edge research directly in response to SE-related questions, ResearchBot facilitates the application of academic insights to practical contexts. Ultimately, it aims to bridge the gap between academia and industry, using research evidence to support learning and decision-making in software development.
Index Terms:
Software Engineering, Q&A, chatbot
I Introduction
Programmers frequently use Q&A platforms like Stack Overflow [^1] https://stackoverflow.com/ for software engineering support and problem-solving[1]. Similarly, research shows developers often rely on Stack Exchange [^2] https://stackexchange.com/ to seek information about various software development processes and problems [2]. However, while these resources provide a wide range of user-generated answers, prior work suggests responders in virtual knowledge sharing communities may lack expertise and skills necessary to respond to a wide variety of SE-related questions [3].
Recently, large language models (LLMs), e.g., ChatGPT, [^3] https://chat.openai.com/ are capable of providing solutions to SE-related tasks due to their ability to provide contextually relevant answers [4, 5]. In addition, prior work suggests developers prefer responses from ChatGPT over Stack Overflow [6]. However, these tools typically draw answers from general web sources, including blogs and informal Q&A sites, rather than scholarly databases. This can limit the reliability and depth of their answers, as ChatGPT’s responses sometimes include broken or irrelevant links and lack connections to academically validated sources. Prior work highlights these limitations in both Stack Overflow [7] and ChatGPT [8], emphasizing the need for novel approaches that align academic research with practical community needs more effectively. While platforms like Stack Overflow provide quick solutions, ResearchBot offers evidence based insights from scholarly research, ensuring higher reliability and depth for critical decisions in software engineering.
Academic research can offer distinct advantages for practitioners, such as evidence-based insights to enhance SE [9, 10, 11]. However, there is a significant gap between academic research and practical programming communities. Despite the rigorous methods and extensive analyses used in SE research [12], findings from scientific insights often fail to reach practitioners seeking solutions [13, 14]. This gap is caused by challenges such as the relevance and credibility of research when applied to the fast-paced field of software development [15, 16] in addition to a lack of accessibility due to paywalls, technical language, and the volume of available publications [17, 18]. Prior work has explored methods to increase adoption of SE research in practice, such as research summaries of empirical SE findings for practitioners, [^4] https://neverworkintheory.org/ ultimately finding these approaches ineffective [17]. Moreover, sources suggest bridging the academia-industry gap is one of the top priorities for the future of AI [19].
@@ I Introduction
To address these challenges, we introduce ResearchBot, a chatbot designed to support and answer developer questions based on academic research. ResearchBot leverages advanced natural language processing (NLP) techniques to automatically understand user questions, and retrieve relevant papers from the CrossRef repository [^5] https://www.crossref.org/—a comprehensive database containing metadata about scholarly articles, conference proceedings, books, and datasets—and provide concise summaries. This work-in-progress paper presents the results of a preliminary evaluation exploring the capabilities of ResearchBot to provide human-like responses to practical SE-related queries. By transforming theoretical knowledge into practical contexts, ResearchBot aims to provide a first step in bridging the gap between academia and industry—facilitating the application of research evidence to real-world SE problems [11].
II Implementation
Figure 1: Pipeline for ResearchBot
@@ II Implementation
The implementation of ResearchBot integrates two main modules: the Retrieval Module and the Question-Answer (QA) Module, leveraging GPT-4’s capabilities of natural language understanding. This system utilizes specific prompts to guide the model in querying the Crossref repository and generating informative responses based on retrieved research papers. For now, ResearchBot is unable to analyze and generate code as it only gives answers to practical questions based on research findings. Figure 1 illustrates the pipeline of ResearchBot, showcasing the flow from query to response generation.
II-A Retrieval Module
@@
The Retrieval Module fetches relevant academic papers based on queries.
Here is the workflow of retrieval module:
  1. 1.
@@
    (e.g., “nlp+natural+language+processing+document+
@@
Here is the prompt to instruct GPT-4 to query the Crossref API:
From now on I will pass in a prompt like "How do I perform document similarity using NLP" and I would like you to curate an output that looks similar to this "nlp+natural+language+processing+document+ similarity". This is what a search query would look like for the given prompt. It uses ‘+’ which separates all the keywords.
The prompt used in this module guides GPT-4 on how to transform a natural language question into a structured query format. This ensures that the search query generated by the model is specific and optimized for retrieving relevant academic papers. Without a clear prompt, the model might struggle to understand the intended format of the query and could produce less accurate results.
II-A1 Paper Abstract Extraction
To retrieve abstracts for ResearchBot, we utilize the CrossRef API [^6] https://api.crossref.org to initialize the query parameters, setting the number of desired results and filtering for entries with abstracts. We then send a request to the API to fetch papers matching the query, sorted by relevance. For a success request, the response data is parsed to extract paper details, including titles, abstracts, URLs, and citation counts. Each abstract is processed using a function (jats_to_plain_text) that converts the JATS-formatted XML to plain text to ensure readability. Finally, a list of relevant papers sorted by citation count in descending order is created for further analysis.If no relevant abstracts are found, ResearchBot returns a message indicating insufficient research data. Future iterations will incorporate full-text analysis to improve accuracy.
II-B QA Module
@@
Here is the workflow of QA module:
  1. 1.
@@
Here is the prompt to instruct GPT-4 to answer user’s question:
From now on I will pass in a Software Engineering question and some relevant research papers as context for the answer. I would like you to answer the question with references made to the contexts (make sure that the focus of the answer is to answer the question, not the context itself). Be sure to also include the URL after the reference to the paper in the text. Use this format to cite in-text (<a href="url" target="_blank">title</a>).If your question is outside computer science, ResearchBot cannot provide an answer as it only supports CS-related queries.
The prompt used in the QA Module provides GPT-4 with clear instructions on how to structure its response. By specifying the focus on answering the user question with reference to the provided context (research paper abstracts), the prompt ensures that the generated response is relevant and informative. LLMs can excel in understanding and generating human-like text, enabling them to process complex queries and produce responses tailored to specific questions [20]. Additionally, the prompt dictates the citation format, enabling proper acknowledgment of the cited sources within the response.
III Methodology
III-A Data Collection
@@
For the purpose of this study, data was systematically collected from the Software Engineering Stack Exchange community. [^7] https://softwareengineering.stackexchange.com/ This platform was selected due to its extensive use by programming professionals, who regularly engage in discussions concerning practical challenges and their solutions. Our selection criteria focused primarily on technical inquiries, particularly those pertaining to subjects such as programming languages and software development. We focused on technical inquiries from SE Stack Exchange, avoiding Stack Overflow due to coding-related limitations. Using these criteria, we collected a dataset of 100 question-answer pairs to analyze the functionality of ResearchBot.
III-B Data Analysis
III-B1 Semantic Textual Similarity
@@
We utilized semantic textual similarity (STS), an NLP technique for determining how similar two pieces of text are in terms of meaning, to analyze our ResearchBot’s performance and accuracy. The primary goal of this analysis is to compare the likeness in meaning between the user’s answer to a question on SE Stack Exchange and ResearchBot’s generated response. The higher their similarity, the better our tool answers the user’s question. To achieve our goal, we used the Sentence-BERT  [^8] https://sbert.net/ (SBERT) model and the similarity method from spaCy  [^9] https://spacy.io/usage/linguistic-features#vectors-similarity, an open-source NLP library in Python.
III-B2 Analysis Process
@@
We selected 100 question-and-answer sets for analysis. We first wrote a script to generate corresponding answer outputs for the 100 questions by ResearchBot. Next, we used the STS tools, SBERT and spaCy, to compare the answers most upvoted from users (top answer) and answers produced by ResearchBot (generated answer). We used the top answer to represent the correct answer, most accepted by humans with similar inquiries. Finally, we calculated each comparison set’s semantic similarity score and each STS tool’s overall statistic measurement characteristics of data. This analysis may introduce biases, as it is based on the assumption that the top answer is both correct and precise. This could lead to overestimating the bot’s performance, as similarity with the top answer might not represent the most accurate or comprehensive response. To enhance the validity of our results, the incorporation of a broader range of benchmarks and metrics are recommended.
III-B3 LLM Comparison
To further assess ResearchBot’s performance, we compared its responses with those generated by other LLM tools commonly used in SE: Google Gemini, GitHub Copilot, and ChatGPT. This evaluation involved the authors manually comparing the 50 top questions from Stack Exchange with responses from the selected LLMs. We specifically looked for the evidence provided to back up responses for each model. Upon manual review of the answers, we found responses from Gemini and Copilot were unreliable and lacked quality references. ChatGPT offered better resources; thus, as part of our future work we plan to conduct a user study in which SE practitioners evaluate responses from both ResearchBot and ChatGPT. This will provide insights into user preferences regarding the clarity, reliability, and overall usefulness of responses, enabling us to refine ResearchBot based on practitioner feedback and further tailor it to meet their needs.
IV Results
@@
This section will present the results of our data collection and analysis approach. Table I presents the statistical results for semantic similarity comparing the performance of ResearchBot with Stack Exchange.
TABLE I: Stack Exchange vs. ResearchBot
@@ IV-A STS Results
The results presented in Table I demonstrate that ResearchBot performs effectively in answering users’ questions as analyzed by spaCy, with an average similarity score of 0.92. However, the performance analyzed by SBERT shows a lower average of 0.66. This discrepancy is largely due to a notably low minimum similarity score of 0.0341 recorded by SBERT. This result highlights SBERT’s higher scrutiny evaluation criteria, contributing to higher variability in scores, as indicated by a standard deviation of 0.16, compared to spaCy. Such variability suggests a re-evaluation of spaCy’s data handling, as it indicates methodological differences, suggesting a need for a deeper comparative analysis of how each tool handles and interprets textual variations. Further analysis could provide insights into the different sensitivities of the tools to specific text features or contexts, thus explaining the variations in their performance metrics. Moreover, similar to recent efforts in sentiment analysis, future work can explore SE-specific NLP tools for STS.
IV-B LLM Comparison
The complete analysis for our manual analysis is available online. [^10] https://shorturl.at/QUAhn Overall, the answer outputs demonstrate ResearchBot provides more reliable responses than other LLMs. For instance, we observed 12% of responses provided by Gemini ($n=6$) contained at least one invalid URL. Across GPT-4o, Gemini, and GitHub Copilot, we observed no references to academic resources in responses, with most relying only on blog posts and other non-credible sources [21]—if evidence was provided at all. Meanwhile, all responses from ResearchBot ($n=50$, 100%) incorporated scholarly literature in generated responses to a variety of SE-related questions.
@@ IV-B LLM Comparison
Our findings show ResearchBot not only matches the top answers but also provides more comprehensive and detailed responses with research-based evidence. This makes our tool particularly valuable. Consequently, while ResearchBot shows promise, there is room for improvements of our tool and evaluating its capabilities in answering SE-related questions.
V Discussion
V-A Implications of the Results
@@
By utilizing STS tools and comparing against mainstream LLMs, we evaluated ResearchBot’s ability to generate accurate and reliable responses. The results show mean STS scores of 0.66 (SBERT) and 0.92 (spaCy), indicating human-like responses to SE-related questions. Despite SBERT’s lower scores, manual analysis confirms that ResearchBot provides not only correct answers but also valuable background information. Our system’s responses are comparable to ChatGPT while offering more reliable, research-backed answers. This study highlights the potential for evidence-based tools to enhance knowledge transfer between academia and industry in SE.
V-B Future Work
@@
User Study to Further Evaluate the Responses
A limitation of our preliminary study is we rely on STS comparisons with top answers from SE Stack Exchange and ChatGPT responses to evaluate the capabilities of ResearchBot. As previously mentioned, we are currently conducting a user study where participants are asked to choose between answers to practical SE-related questions generated by ResearchBot and ChatGPT, providing a rationale for their preference. This study aims to provide a better understanding of the types of answers users prefer, motivate enhancements to ResearchBot, and further promote evidence-based practices in SE contexts. Future studies will include usability testing with software developers to assess the chatbot’s interface, response clarity, and practical relevance.
Expansion of Academic Repositories
@@
User Interaction and Feedback Mechanisms
@@
Code Analysis
@@
Integration with Programming Platforms
We also envision exploring integration opportunities for our system in popular programming platforms beyond SE Stack Exchange, such as GitHub or developer forums. This can expand ResearchBot’s reach and utility within the broader SE community. Further, integration into integrated development environments (IDEs) can also facilitate insights from ResearchBot during software development to answer questions and inform decision-making based on empirical research findings.
VI Conclusion